Alex Yurechuk 1878-1963  Mary Antoniuk 1882-1978 m Stephen Yurechuk 1902-1971 Katherine Kolody 1908-  Stephen Michajluk 1894-1977 Donna Yurechuk 1905-1991  William Finnie  Ellen Yurechuk 1907-  William Kassian 1905-  Anna Yurechuk 1910-1989  William Yurechuk 1913-1989  Olga Maksymic 1923-1976  David Bengtson 1914-  Lillian Yurechuk 1915-1966  William (Formanski) Forman 1913-1986 William (Simkovich) Simcoe  Mary Yurechuk 1917-1972  Alexander Yurechuk 1924-  Gladys Lundeby 1924- m  m 1927  m  m  m  m 1941?  m  m  m Andrew (Yurechuk) Stephen 1927-  Anne ?  Elizabeth Josephine Summers Terrance (Yurechuk) Stephen 1930-  Stella Piasecki 1930-  Shirley A. Madison 1934- m  m 1957  m 1951  m 1981 Jane Elizabeth Stephen 1958-  Boyd Andrew  Ross Terrence Stephen 1956-  Dianne Schmitz  Rodney A. Stephen 1959-  Cheryl Caper  Stacey-Jo Stephen 1963- m 1982  m Tyler Stephen  Scott Stephen  Heather Stephen 1991-  Ernest George Lake 1928-  Mary Pauline Michajluk 1931-  William Stephen Michajluk 1936-  Ann Elizabeth Dagg m 1952  m 1964 Randall George Lake 1954-  Robert Murray Lake 1957-  Ronald John Willis 1960-  Debra Marilyn Lake 1959- m 1985 Shannon Marie Willis 1988-  Alanna Michelle Willis 1989-  Shawn Douglas Michajluk 1967-  Glen Robertson  Lisa Suzanne Michajluk 1970-  Peter William Michajluk 1972- m 1996  Terence Kassian 1933-  Wilma Steckel  Leonard Kassian  Mabel Yewchuk m  m Robert Kassian  Wayne Kassian  Carol Kassian  Brian Kassian  Lisa Kassian  William Yurechuk 1942-  Gladys Skordbahach  Douglas Kempf  Elizabeth Yurechuk 1944-  Douglas Pushor  Patricia Yurechuk 1955-  Terrence Fosvelt  Sandra Yurechuk 1961- m  m  m  m Carrie Yurechuk Stephen Yurechuk  Stacey Yurechuk  Michael Yurechuk  Fred Kempf  Jo Kempf  Sheila Pushor  Bonnie Pushor  Colleen Pushor  Brandi Fosvelt  Dustin Fosvelt  Roald Dyrland 1939-  Judith Bengtson 1943-  Clifton Bengtson 1944-  Kay Luna m 1963  m  Gerald Forman 1937-  Rose ?  Vivian ?  Robert Forman 1939-  Patricia Mallick 1941-1993  Kenneth Forman 1942-1992  Sherry ? m  m  m  m Son  Son  Son  Daughter  Darren Forman  Kathy Forman  Son  Gerald Vetsch 1951-  Shirley Yurechuk 1951-  Gerald Kewchuk 1955-  Laurie Yurechuk 1955- m  m Shauna Vetsch 1974-  Ryan Vetsch 1976-  Brandon Vetsch 1978-  Craig Vetsch 1981-  Bradley Kewchuk 1983-  Jennifer Kewchuk 1987-

1